Circular No. 7090

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999E IN ANONYMOUS GALAXY A. V. Filippenko, D. C. Leonard, and A. G. Riess, University of Cal-

ifornia at Berkeley, report: “A spectrum (range 510–880 nm) obtained on Jan. 19 UT with the Keck-II telescope reveals that SN 1999E (cf. IAUC 7089) is a very unusual type-IIn supernova at redshift 0.025. Besides the strong, relatively narrow Hα emission line on top of a broader base, there are very broad (FWHM ∼ 40 nm) emission bumps centered near 560, 640, 755, and 875 nm. The spectrum differs from that of any other supernova ever examined by the authors.”

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, write: “A spectrum of SN 1999E, obtained by P. Berlind on Jan. 19.5 UT with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph), displays strong, narrow (FWHM 370 km/s) Hα emission superimposed on a broad (FWHM 7000 km/s) base. Weak emission from Hβ is also present. A strong interstellar Na I D absorption feature (equivalent width 0.1 nm) from the host galaxy provides a recession velocity of 7800 km/s. These features suggest that the supernova is of type IIn, and significantly reddened. One peculiarity is a broad (FWHM 18 000 km/s), unidentified absorption feature observed at 597 nm, not typically seen in type-IIn supernovae. The supernova is quite luminous; for V = 16 as reported by Perez et al. (IAUC 7089) and H_{0} = 65 km s^{−1} Mpc^{−1}, the absolute magnitude is M_{V} = −19.4. Given the reddening, the intrinsic luminosity of the supernova is likely significantly brighter.”

SUPERNOVA 1998fa IN UGC 3513 Filippenko et al. also report that a spectrum (range 510–880 nm) ob-

tained on Jan. 19 UT with the Keck-II telescope, in comparison with previous spectra, reveals that SN 1998fa (cf. IAUC 7073, 7077) is actually an unusual ‘type-IIb’ supernova like SN 1987K (Filippenko 1988, A.J. 96, 1941) and SN 1993J (Filippenko et al. 1993, Ap.J. 415, L103): lines of He I with P-Cyg profiles have emerged.

SUPERNOVA 1998ez IN ANONYMOUS GALAXY Filippenko et al. further report that a spectrum obtained as above on

Jan. 20 UT reveals that SN 1998ez (cf. IAUC 7070) is of type II at redshift 0.07; Hα emission is visible on a mostly featureless continuum.

1999 January 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green